A first order reaction is 30 percent completed in 35 minute what is the value of the rate constant?

b) whay percent of the reactant will remain after 5 hours?


1
Expert's answer
2021-05-12T06:27:19-0400

Solution:

Let 1 M - the initial concentration, [A]o.

After 35 minutes, the concentration, [A], is: 1 - 0.3 = 0.7 (M)


a)

For a first order reaction:

ln[A] = −kt + ln[A]o

Hence,

ln(0.7) = −k × 35 + ln(1)

ln(0.7) = −k × 35

-0.3567 = −k × 35

k = 0.01019 min-1


b)

(5 h) × (60 min / 1 h) = 300 min

ln[A] = −kt + ln[A]o

ln[A] = (−0.01019 × 300) + ln(1)

ln[A] = -3.057

[A] = e-3.057 = 0.047

0.047 M × 100% / 1 M = 4.7%

So, 4.7% of the reactant will remain after 5 hours.


Answers:

a) The value of the rate constant (k) is 0.01019 min-1

b) 4.7% of the reactant will remain after 5 hours